OpenStudy (anonymous):

Find the quotient. (5ab - 10b2 + 15bc) ÷ 5b a + 2b - 3c a - 2b + 3c a - 2b - 3c

Parth (parthkohli):

\(\Large \color{MidnightBlue}{\Rightarrow 5b(?) = 5ab - 10b^2 + 15bc }\) Question mark will be the answer xD

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

?

Parth (parthkohli):

What do you multiply to 5b to get 5ab - 10b^2 + 15bc?

OpenStudy (anonymous):

2

OpenStudy (merchandize):

= (5ab - 10b² + 15bc)/5b = (5b[a - 2b + 3c])/5b = a - 2b + 3c Answer: a - 2b + 3c

OpenStudy (anonymous):

thanks